OpenStudy (anonymous):

At what rate of interest per annum will an amount get doubled in 16 years? A- 12% B- 8% C- 15% D- 25/4%

OpenStudy (wolf1728):

Here's a formula log(1 + rate) = {log(total) -log(Principal)} ÷ Years Don't be intimidated. Let's say the total is 2 and the principal is 1 and years = 16 log (1+rate) = log(2) - log(1) / 16 log(1 + rate) = (0.3010299957 - 0) /16 log(1 + rate) = 0.0188143747 10^0.0188143747 = 1 + rate 1.0442737824 = 1 + rate .0442737824 = rate or rate = 4.42737824

OpenStudy (wolf1728):

I got the formula from here: http://www.1728.org/compint2.htm And to check that answer, there's a formula and a calculator here: http://www.1728.org/compint.htm Total = Principal * (1 + rate) ^ years 2 = 1 * (1.0442737824)^16 2 = 2 It checks

OpenStudy (anonymous):

but 2 is not in their options

OpenStudy (anonymous):

which one is right?

OpenStudy (wolf1728):

No 2 is not the answer. The RATE is 4.42737824 which is close to D 25/4

OpenStudy (wolf1728):

Wait that comes to 6.25 % Let me recheck that

OpenStudy (anonymous):

http://skillgun.com/question/4814/arithmetic/simple-interest/at-what-rate-percent-per-annum-will-a-sum-of-money-double-in-16-years in this link have similar question but ans process is not in their

OpenStudy (wolf1728):

The formula for compound interest is Total = principal * (1+rate)^years we need the total to be twice what we started with so let's say total = 6 and principal =3 6 = 3 * (1.0442737824)^16 6 = 5.9999999975 Close enough don't you think?

OpenStudy (anonymous):

but which one i choose the option

OpenStudy (wolf1728):

The formula they use isn't the correct one

OpenStudy (wolf1728):

Do you have access to an online compound interest calculator so you can try 6.25% to see that it is INCORRECT?

OpenStudy (anonymous):

ya i was try thats incorrect?

OpenStudy (wolf1728):

I found one here: http://www.moneychimp.com/calculator/compound_interest_calculator.htm My answer 200 at 6.25% interst produces an answer of 527.59 That should be $400 but it isn't

OpenStudy (wolf1728):

I went to that calculator and used $200 for 16 years at 4.42737824% and my answer was $400.16 Seems as if I'm right. I could find another calculator.

OpenStudy (wolf1728):

Here's another online calculator http://www.calculatorsoup.com/calculators/financial/compound-interest-calculator.php I input $100 4.42737824% rate and 16 years answer was $200.00

OpenStudy (irishboy123):

answer is D **simple** interest at rate r for 16 years on amount X equals 16 * r * X set that to X to double the original amount ie solve 16 * r * X = X =>> r = 1/16 = 100/16 % = 25/4% the 4.43% is if interest is compounded in which case you are solving this for r

OpenStudy (irishboy123):

this! \[X (1+r)^{16} = 2X\]

OpenStudy (wolf1728):

Where are you getting that formula?

OpenStudy (irishboy123):

which formula?? if you mean for **compounded** interest , it is just this start amount X after period 1 X(1+r) after period 2 X(1+r)(1+r) = X (1+r)^2 after period 3 X(1+r)(1+r)(1+r) = X (1+r)^3 and so on it's really that simple.

OpenStudy (wolf1728):

You mean this is SIMPLE interest? Geez, I'm so used to solving compound interest problems. (They are a "tad" more difficult).

OpenStudy (irishboy123):

but the question is about simple interest so you just add up all the interest as its earned, but you never add it to the interest bearing original amount so 16 * r * X = X

OpenStudy (irishboy123):

@wolf1728 yes mate it should have been stated in the question. your solution is sound but addresses a more complex question!

OpenStudy (wolf1728):

Yes, I understand simple interest. This is the FIRST time I even saw a simple interest question here.

OpenStudy (anonymous):

yes thanks to all

OpenStudy (wolf1728):

okay